Grade 12th passMechanics

a certain force gives object mass m1 an acceleration of 12 and mass m2 an acceleration of 3.8 m/s2. What acceleration would the force give to an object with a mass of (m2-m1)? What acceleration would the force give to an object with a mass of (m2+m1)?

zubaria
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear student
 
1) m1 = F/12
2) m2 = F/3.8
Combining, m2/m1 = 3.1578
Ergo, m2-m1 = m1*(3.1578- 1) = 2.1578m1 

a(2-1) = 12/2.1578 = 5.56 m/s² 

a(2+1) = 12/(3.1578+1) = 2.886 m/s²
 
Regards
Arun (askIITians forum expert)